Problem 1 Let X 1 , X 2 , ⋯,Xn be a random sample from a normal distribution with mean µ and variance σ 2 . Consider e X as an estimator of e µ where X is the sample mean. Show that e is consistent estimator of e µ X .


1
Expert's answer
2021-06-07T07:32:02-0400

"E(estimator)=E(\frac{1}{n}\displaystyle\sum_{i=1}^nX)"


"E(\bar{X})=\frac{1}{n}\sum E(xi)"


"E(\bar{X})=\frac{1}{n}\displaystyle\sum_{i=1}^nu"


"=E(\bar{X})=\frac{1}{n}.nu"


"E(\bar{X})=u"